A machinist wishes to insert a brass rod with a diameter of 2 mm into a hole with a diameter of 1.995 mm. By how much would the machinist have to lower the temperature (in °C) of the rod to make it fit the hole?


1
Expert's answer
2022-02-13T12:15:56-0500

"\Delta T=\frac{\Delta l}{\alpha l}=\frac{5\cdot 10^{-6}}{19\cdot 10^{-6}\cdot 2\cdot 10^{-3}}=132°C."